# Generalized Out-of-Distribution Detection and Beyond in Vision Language Model Era: A Survey

### v1

@@ Abstract
Detecting out-of-distribution (OOD) samples is crucial for ensuring the safety of machine learning systems and has shaped the field of OOD detection. Meanwhile, several other problems are closely related to OOD detection, including anomaly detection (AD), novelty detection (ND), open set recognition (OSR), and outlier detection (OD). To unify these problems, a generalized OOD detection framework was proposed, taxonomically categorizing these five problems. However, Vision Language Models (VLMs) such as CLIP have significantly changed the paradigm and blurred the boundaries between these fields, again confusing researchers. In this survey, we first present a generalized OOD detection v2, encapsulating the evolution of AD, ND, OSR, OOD detection, and OD in the VLM era. Our framework reveals that, with some field inactivity and integration, the demanding challenges have become OOD detection and AD. In addition, we also highlight the significant shift in the definition, problem settings, and benchmarks; we thus feature a comprehensive review of the methodology for OOD detection, including the discussion over other related tasks to clarify their relationship to OOD detection. Finally, we explore the advancements in the emerging Large Vision Language Model (LVLM) era, such as GPT-4V. We conclude this survey with open challenges and future directions.
Index Terms:
Anomaly Detection, Novelty Detection, Open Set Recognition, Out-of-Distribution Detection, Outlier Detection, Vision Language Model, CLIP, Large Vision Language Model, Large Multi-modal Model
@@ 1 Introduction
Areliable visual recognition system should not only accurately predict known contexts, but also identify and reject unknown examples [1, 2, 3, 4]. In critical applications like autonomous driving, the system must alert and cede control to the driver upon encountering unfamiliar scenes or objects not seen during training. However, most existing machine learning models are trained based on the closed-world assumption [5, 6], where the test data is assumed to be drawn i.i.d. from the same distribution as the training data, known as in-distribution (ID). Therefore, the development of classifiers capable of detecting out-of-distribution (OOD) samples is a crucial challenge for real-world applications. This challenge is precisely the focus of research in the field of OOD detection.
@@
While OOD detection primarily focuses on semantic distribution shift, several other tasks share similar goals and motivations, including outlier detection (OD) [7, 8, 9, 10], anomaly detection (AD)  [11, 12, 13, 14], novelty detection (ND)  [15, 16, 17, 18], and open set recognition (OSR) [19, 20, 21]. Subtle differences in the specific definitions among these sub-topics have caused confusion in the field, leading to similar approaches being proposed across them. To address this issue, the generalized OOD detection framework was introduced [22]. The taxonomy of the generalized OOD detection framework is shown in Fig. 1. The generalized OOD detection framework classifies these tasks as special cases or sub-tasks under a unified taxonomy. This framework provides clear definitions and fosters a deeper understanding of each field.
Figure 1: Taxonomy of generalized OOD detection framework [22], illustrated by classification tasks. Figure adapted from [22].
@@
In recent years, the emergence of Vision Language Models (VLMs), represented by CLIP [23], has rapidly accelerated research in the field of Computer Vision. This has changed the paradigm of the recognition field, allowing for zero-shot [23] or few-shot learning [24, 25] in various domains. VLMs have significantly influenced the aforementioned five problems (OD, AD, ND, OSR, and OOD detection), and the application of VLMs has become a highly notable research field [26, 27, 28, 29]. However, alongside this remarkable progress, the paradigm shift with the advent of the VLMs has blurred the boundaries between the five problems. Due to the difficulty of a clear understanding of the distinctions and interrelations between these tasks, each community within the fields is facing significant challenges in identifying the optimal direction to pursue in this VLM era.
@@
Finally, we introduce the evolution of these problems in the emerging Large Vision Language Model (LVLM) era, such as GPT-4V [30] or LLaVA [31] (also referred to as Large Multi-modal Models or Multi-modal Large Language Models [32]). We summarize the definition of each evolving problem, the findings so far, and future challenges.
TABLE I: Number of VLM-based papers in the Top Venues by June 2024.
Task Top Venue
(a) Sensory AD CVPR2023$\times$[27] ICLR2024$\times$[29] CVPR2024$\times$[33, 34, 35, 36]
(b) Semantic AD/ND TMLR2022$\times$[37] (one-class) CVPR2024$\times$[35] (one/multi-class)
(c) OSR None
(d) OOD Detection NeurIPS2021$\times$[38] AAAI2022$\times$[39] NeurIPS2022$\times$[26] ICCV2023$\times$[40] IJCV2023$\times$[41] NeurIPS2023$\times$[28, 42, 43] ICLR2024$\times$[44, 45] CVPR2024$\times$[46, 47] ICML2024$\times$[48]
(e) OD ICML2024 [49]
To sum up, in this survey paper, we make three contributions to the research community:
@@
    Unified Framework in the VLM era: We systematically review the evolution of five closely related topics of OD, AD, ND, OSR, and OOD detection in the Vision Language Model era and provide an updated unified framework termed generalized OOD detection v2. Our framework reveals that the paradigm shift has led to some field inactivity or integration, and the demanding challenges are AD and OOD detection. We hope that these observations highlight the demanding challenges in the VLM era and foster collaborative efforts among each community.
@@
    A Comprehensive Survey for OOD Detection in the VLM era: While comprehensive surveys on OD, AD, ND, OSR, and OOD detection methodologies have been published in recent years [11, 12, 13, 14, 20, 22, 50, 51], this survey is the first to comprehensively overview OOD detection methods specifically in the VLM era. By connecting with other related tasks, we aim to provide readers with a holistic understanding of the developments and interconnections within each problem, particularly in the context of OOD detection.
@@
    An Introduction to the Evolution in the LVLM Era: We further introduce the evolution of each problem in the Large Vision Language Model era. Despite the infant stage of these fields, this survey offers an in-depth introduction to each problem, aiming to facilitate future advancements in this area.
@@
    Future Research Directions: We draw readers’ attention to the future work necessary for advancing the field in the VLM and LVLM era. We conclude this survey with discussions on open challenges and opportunities for future research.
Figure 2: Generalized OOD detection framework v2, reflecting the evolution of each problem in the VLM era. (a) Sensory AD has become a highly active and noteworthy field in the VLM era. In terms of benchmarks, in addition to the commonly used MVTec-AD [52], VisA [53], the largest industrial anomaly detection dataset, has also become a standard benchmark in the field. (b) Semantic AD/ND has become inactive in the VLM era. (c) OSR has been integrated into hard OOD detection. CLIP-based hard OOD detection incorporates the benchmark setup of OSR and creates new benchmarks such as ImageNet-10/ImageNet-20 [26] and ImageNet-protocol [54, 47]. (d) OOD detection is a highly active research area in the VLM era. (e) OD has become inactive in the VLM era.
The paper content is organized as follows. In Sec. 2, we introduce the new version of generalized OOD detection by summarizing the evolution of the five related fields in the VLM era. We then overview the two key problems (OOD detection and AD) that have evolved and remain active in Sec. 3, with a detailed breakdown of existing methodologies being presented in Sec. 4 (CLIP-based OOD detection) and Sec. 5 (CLIP-based AD). In Sec. 6, we introduce early advancements of OOD detection and AD in the LVLM era. Sec. 7 and Sec. 8 feature discussions on potential challenges and future directions. Finally, we conclude with Sec. 9.
@@ 2 Generalized OOD Detection V2
In this section, we introduce a novel unified framework termed generalized OOD detection v2, which summarizes the evolution of the five related fields in the VLM era. We first revisit the previous generalized OOD detection framework in Sec. 2.1. Next, we introduce the evolution of each problem.
@@ 2.1 Background: Generalized OOD Detection V1
To provide a clear definition, a generalized OOD detection framework was proposed [22]. The taxonomy for generalized OOD detection is shown in Fig. 1. It is based on the following four bases: (1) Distribution shift to detect: The task focuses on detecting either covariate shift (e.g., OOD samples from a different domain) or semantic shift (e.g., OOD samples from a different semantic). (2) ID data type: The in-distribution (ID) data contains either a single class or multiple classes. (3) Whether the task requires ID classification: Some tasks require classification of the ID data, while others do not. (4) Transductive vs. inductive learning: Transductive tasks require all observations (both ID and OOD), while inductive tasks follow the common train-test scheme. According to the above taxonomy, these five problems can be clearly categorized as shown in Fig. 1: Anomaly detection is categorized into sensory anomaly detection, which deals with covariate shift, and semantic anomaly detection, which deals with semantic shift. Novelty detection falls under the same category as semantic anomaly detection. When addressing a multi-class scenario that necessitates ID classification, both open-set recognition and out-of-distribution detection are encompassed within this category. The main difference between OSR and OOD detection was the benchmark setup [22, 55] (Sec. 2.2 (c)). Outlier detection belongs to a different category from the other tasks, as this problem is transductive (i.e., it has access to all observations).
For the detailed definition of each task, we refer the readers to the previous generalized OOD detection survey paper [22].
2.2 Evolution of Each Problem in VLM Era
We review how each problem has evolved in the VLM era. To make a fair judgment, we comprehensively investigated papers that use VLMs from top venues and summarized them in Table I. Our survey revealed that CLIP [23] is predominantly used as the VLM for OOD detection and other sub-tasks, and other VLMs [56, 57] are rarely utilized. Therefore, we focus on CLIP as the target VLM in this survey and refer to OOD detection using CLIP as CLIP-based OOD detection. Similarly, we will prefix other tasks with “CLIP-based” (e.g., CLIP-based AD). As OOD detection research is primarily focused on the image domain, we conduct a survey of other tasks within the image domain that are common and have strong connections to OOD detection research. For instance, our survey does not cover video domain tasks [58, 59, 60] due to their limited connection to OOD detection.

### v2

@@ Abstract
Detecting out-of-distribution (OOD) samples is crucial for ensuring the safety of machine learning systems and has shaped the field of OOD detection. Meanwhile, several other problems are closely related to OOD detection, including anomaly detection (AD), novelty detection (ND), open set recognition (OSR), and outlier detection (OD). To unify these problems, a generalized OOD detection framework was proposed, taxonomically categorizing these five problems. However, Vision Language Models (VLMs) such as CLIP have significantly changed the paradigm and blurred the boundaries between these fields, again confusing researchers. In this survey, we first present a generalized OOD detection v2, encapsulating the evolution of these fields in the VLM era. Our framework reveals that, with some field inactivity and integration, the demanding challenges have become OOD detection and AD. Then, we highlight the significant shift in the definition, problem settings, and benchmarks; we thus feature a comprehensive review of the methodology for OOD detection and related tasks to clarify their relationship to OOD detection. Finally, we explore the advancements in the emerging Large Vision Language Model (LVLM) era, such as GPT-4V. We conclude with open challenges and future directions. The resource is available at https://github.com/AtsuMiyai/Awesome-OOD-VLM.
@@ 1 Introduction
A reliable visual recognition system should not only accurately predict known contexts, but also identify and reject unknown examples (Amodei et al., 2016; Mohseni et al., 2021; Hendrycks et al., 2021b; Hendrycks & Mazeika, 2022). In critical applications such as autonomous driving, the system must alert and cede control to the driver upon encountering unfamiliar scenes or objects not seen during training. However, most existing machine learning models are trained based on the closed-world assumption (Krizhevsky et al., 2012; He et al., 2015), where the test data is assumed to be drawn i.i.d. from the same distribution as the training data, known as in-distribution (ID). Therefore, the development of classifiers capable of detecting out-of-distribution (OOD) samples is a crucial challenge for real-world applications. This challenge is precisely the focus of research in the field of OOD detection.
@@
While OOD detection primarily focuses on semantic distribution shift, several other tasks share similar goals and motivations, including outlier detection (OD) (Aggarwal & Yu, 2001; Hodge & Austin, 2004; Ben-Gal, 2005; Wang et al., 2019a), anomaly detection (AD)  (Ruff et al., 2021; Pang et al., 2021; Bulusu et al., 2020; Chalapathy & Chawla, 2019), novelty detection (ND)  (Pimentel et al., 2014; Miljković, 2010; Markou & Singh, 2003a; b), and open set recognition (OSR) (Boult et al., 2019; Geng et al., 2020; Mahdavi & Carvalho, 2021). Subtle differences in the specific definitions among these sub-topics have caused confusion in the field, leading to similar approaches being proposed across them.
To address this issue, the generalized OOD detection framework was introduced (Yang et al., 2024). The taxonomy of the generalized OOD detection framework is shown in Fig. 1. The generalized OOD detection framework introduces a taxonomy built upon four criteria: distribution shift type (covariate/semantic shift), the type of ID data (single/multi-class), necessity of ID classification, and learning setting (transductive/inductive). According to the above taxonomy, these five problems can be clearly categorized as shown in Fig. 1: AD is categorized into sensory AD, which deals with covariate shift, and semantic AD, which deals with semantic shift. ND falls under the same category as semantic AD. In multi-class settings requiring ID classification, both OSR and OOD detection are included. OD belongs to a transductive category (i.e., it has access to all observations). This framework provides clear definitions and fosters a deeper understanding of each field.
Figure 1: Taxonomy of generalized OOD detection framework (Yang et al., 2024), illustrated by classification tasks. (a) Sensory anomaly detection is categorized as covariate shift detection. (b) Semantic anomaly detection and novelty detection fall under semantic shift detection. (c) Open set recognition and (d) out-of-distribution detection are classified as multi-class semantic shift detection tasks that require ID classification. (e) Outlier detection is characterized by having a transductive observation type.
@@
In recent years, the emergence of Vision Language Models (VLMs), represented by CLIP (Radford et al., 2021), has rapidly accelerated research in the field of computer vision. This has changed the paradigm of the recognition field, allowing for zero-shot (Radford et al., 2021) or few-shot learning (Zhou et al., 2022c; b) in various domains. VLMs have significantly influenced the aforementioned five problems (OD, AD, ND, OSR, and OOD detection), and the application of VLMs, particularly CLIP, has become a highly notable research field (Ming et al., 2022a; Jeong et al., 2023; Miyai et al., 2023b; Zhou et al., 2024a). However, alongside this remarkable progress, the paradigm shift with the advent of the VLMs has blurred the boundaries between the five problems. Due to the difficulty of a clear understanding of the distinctions and interrelations between these tasks, each community within the fields is facing significant challenges in identifying the optimal direction to pursue in this VLM era.
@@
Finally, we introduce the early evolution of these problems in the emerging Large Vision Language Model (LVLM) era, such as GPT-4V (OpenAI et al., 2023) or LLaVA (Liu et al., 2023a; 2024b). We summarize the definition of each evolving problem, the findings so far, and future challenges.
Table 1: Number of VLM-based papers in the Top Venues from 2021 to April 2025
. Task Top Venue (a) Sensory AD CVPR2023$\times$(Jeong et al., 2023), ICLR2024$\times$(Zhou et al., 2024a), CVPR2024$\times$(Li et al., 2024c; Ho et al., 2024; Zhu & Pang, 2024; Huang et al., 2024), ACMMM×$\times$(Gu et al., 2024a; Zhu et al., 2024a), IJCAI2024$\times$(Zuo et al., 2024), ECCV2024$\times$(Cao et al., 2024c; Qu et al., 2024), NeurIPS2024$\times$(Li et al., 2024f; Arodi et al., 2024), ICLR2025$\times$(Jiang et al., 2025; Yun et al., 2025), CVPR2025$\times$(Sun et al., 2025; Ma et al., 2025; Qu et al., 2025; Zhang et al., 2025b; Fan et al., 2025; Gu et al., 2025) (b) Semantic AD/ND TMLR2022$\times$(Liznerski et al., 2022), CVPR2024$\times$(Zhu & Pang, 2024), ICLR2025$\times$(Yun et al., 2025) (c) OSR ECCV2024$\times$(Miller et al., 2024) (d) OOD Detection NeurIPS2021$\times$(Fort et al., 2021), AAAI2022$\times$(Esmaeilpour et al., 2022), NeurIPS2022$\times$(Ming et al., 2022a), ICCV2023$\times$(Wang et al., 2023a), IJCV2023$\times$(Ming & Li, 2024a), NeurIPS2023$\times$(Miyai et al., 2023b; Tu et al., 2023; Park et al., 2023; Liu et al., 2023b), ICLR2024$\times$(Nie et al., 2023; Jiang et al., 2024), CVPR2024$\times$(Bai et al., 2024a; Li et al., 2024b), ICML2024$\times$(Cao et al., 2024a), ECCV2024$\times$(Zhang et al., 2024d; Liu & Christopher, 2024; Zhang et al., 2024f; Lafon et al., 2024), NeurIPS2024$\times$(Li et al., 2024e; Yu et al., 2024; Chen et al., 2024a; Zhang et al., 2024a; Zhang & Zhang, 2024), TMLR2024$\times$(Adaloglou et al., 2024), IJCV2025$\times$(Miyai et al., 2025b), ICLR2025$\times$(Zeng et al., 2025) (e) OD ICML2024$\times$(Liang et al., 2024)
To summarize, this survey presents four contributions to the research field:
@@
    Proposing a Generalized OOD Detection v2: We analyze the progression of five related topics (AD, ND, OSR, OOD detection, and OD) in the VLM era and propose generalized OOD detection v2. Our framework reveals that the paradigm shift has led to some field inactivity or integration, and the demanding challenges are AD and OOD detection. We hope that these observations highlight the demanding challenges in the VLM era and foster collaborative efforts among each community.
@@
    An Extensive Survey for VLM-based OOD Detection: We provide a comprehensive survey of VLM-based OOD detection and AD methods. In particular, we examine recent advances in zero-shot and few-shot settings by categorizing methods according to their training strategies and the use of additional prompts. Although various surveys on OD, AD, ND, OSR, and OOD detection have been conducted (Ruff et al., 2021; Pang et al., 2021; Bulusu et al., 2020; Chalapathy & Chawla, 2019; Geng et al., 2020; Yang et al., 2024; Cao et al., 2024b; Liu et al., 2024c; Xu & Ding, 2025), this work is the first to comprehensively review VLM-based OOD detection methods. By examining the connections among related tasks, we aim to provide insights that enhance the understanding of OOD detection.
@@
    An Introduction to the Evolution in the LVLM Era: We further introduce the evolution of each problem in the LVLM era. Despite the infant stage of these fields, this survey offers an in-depth introduction to each problem, aiming to facilitate future advancements in this area.
@@
    Open Challenges and Future Directions: Finally, we discuss open challenges and future research directions. In particular, by conducting a comparative analysis between the fields of AD and OOD, we identify key areas that are especially important for advancing VLM-based OOD detection. We hope that this survey will serve as a valuable reference for future research on OOD detection and related tasks in the VLM era.
Figure 2: Generalized OOD detection framework v2, reflecting the evolution of each problem in the VLM era. (a) Sensory AD has consistently been an active research field even after the emergence of VLMs. In terms of benchmarks, in addition to the commonly used MVTec-AD (Bergmann et al., 2019), VisA (Zou et al., 2022), the largest industrial anomaly detection dataset, has also become a standard benchmark in the field. (b) Semantic AD/ND has become inactive in the VLM era. (c) OSR has been integrated into hard OOD detection. VLM-based hard OOD detection incorporates the benchmark setup of OSR and creates new benchmarks such as ImageNet-10/ImageNet-20 (Ming et al., 2022a) and ImageNet-protocol (Palechor et al., 2023; Li et al., 2024b). (d) OOD detection is a highly active research area in the VLM era. (e) OD has become inactive in the VLM era.
In the VLM era, research on open-vocabulary segmentation (Gu et al., 2022) and referring segmentation (Wang et al., 2022b) has been related to the discovery of unseen classes. However, these fields mainly focus on tasks that aim to generalize to examples from unseen classes during training when the class name is already known. In contrast, OOD detection and related fields deal with entirely new data, where even the class name is unknown. The goal of OOD detection and related fields is to detect them instead of generalizing to them. Therefore, due to this fundamental difference in motivation and objective, we put the tasks such as open-vocabulary segmentation outside the scope of this survey.
The paper content is organized as follows. In Sec. 2, we introduce the new version of generalized OOD detection by summarizing the evolution of the five related fields in the VLM era. We then overview the two key problems (OOD detection and AD) that have evolved and remain active in Sec. 3, with a detailed breakdown of existing methodologies being presented in Sec. 4 (VLM-based OOD detection) and Sec. 5 (VLM-based AD). In Sec. 7, we introduce early advancements of OOD detection and AD in the LVLM era. Sec. 8 features future directions. Finally, we conclude with Sec. 9.
@@ 2 Generalized OOD Detection V2
In this section, we introduce a novel unified framework termed generalized OOD detection v2, which summarizes the evolution of the five related fields in the VLM era. We first revisit the previous generalized OOD detection framework in Sec. 2.1. Next, in Sec. 2.2, we show the results of the investigation of the research activity in each field. In Sec. 2.3, we introduce the evolution of each problem. Finally, in Sec. 2.4, we have a discussion about the future directions.
@@ 2.1 Background: Generalized OOD Detection V1
To provide a clear definition, a generalized OOD detection framework was proposed (Yang et al., 2024). The taxonomy for generalized OOD detection is shown in Fig. 1. It is based on the following four bases: (1) Distribution shift to detect: The task focuses on detecting either covariate shift (e.g., OOD samples from a different domain) or semantic shift (e.g., OOD samples from a different semantic). (2) ID data type: The in-distribution (ID) data contains either a single class or multiple classes. (3) Whether the task requires ID classification: Some tasks require classification of the ID data, while others do not. (4) Transductive vs. inductive learning: Transductive tasks require all observations (both ID and OOD), while inductive tasks follow the common train-test scheme. According to the above taxonomy, these five problems can be clearly categorized as shown in Fig. 1: Anomaly detection is categorized into sensory anomaly detection, which deals with covariate shift, and semantic anomaly detection, which deals with semantic shift. Novelty detection falls under the same category as semantic anomaly detection. When addressing a multi-class scenario that necessitates ID classification, both open-set recognition and out-of-distribution detection are encompassed within this category. The main difference between OSR and OOD detection was the benchmark setup (Yang et al., 2024; Salehi et al., 2022) (Sec. 2.3 (c)). Outlier detection belongs to a different category from the other tasks, as this problem is transductive (i.e., it has access to all observations).
The more precise definitions of these tasks are as follows:
Anomaly Detection  Anomaly detection (AD) focuses on identifying anomalous instances during inference that deviate from a predefined notion of normality (Chandola et al., 2009). These anomalies can arise from two distinct types of distributional changes: covariate shift and semantic shift (Ruff et al., 2021). Accordingly, AD can be categorized into two primary sub-problems: sensory AD and semantic AD. Sensory AD targets anomalies caused by covariate shift, under the assumption that all normal instances are drawn from the same distribution. In this setting, no semantic shift is considered. In contrast, semantic AD addresses label shift, where normalities are assumed to come from a fixed semantic category. The objective in this setting is to detect test samples that belong to novel, previously unseen classes.
Novelty Detection  Novelty detection (ND) aims to identify test samples that do not belong to any of the categories seen during training. Depending on the number of training classes, ND can be divided into two settings: One-class novelty detection, where the training set contains only one class, and multi-class novelty detection, where multiple classes are present in the training data. It is important to note that, even when multiple ID classes are available, the goal of multi-class novelty detection is solely to distinguish novel samples from known ones, without requiring classification among the ID classes. Both one-class and multi-class ND are binary classification tasks.
Open Set Recognition  Open set recognition (OSR) requires a multi-class classifier to perform two tasks at the same time: (1) correctly classify test samples that belong to known classes, and (2) detect test samples that come from unknown classes. Here, it is common practice to refer to OOD classes as unknown classes and ID classes as known classes. However, there is no difference in meaning between these terms.
Out-of-Distribution Detection  Out-of-distribution (OOD) detection aims to identify test samples that come from a distribution different from the training distribution. In most machine learning tasks, it typically refers to the label distribution, meaning that OOD samples should have labels not seen in the training data. It is also important to note that the training set usually includes multiple classes, and OOD detection should not degrade the model’s ability to classify ID samples correctly.
Outlier Detection  Outlier detection (OD) aims to identify samples that significantly differ from the rest of the given observation set. These differences may arise from either covariate shift or semantic shift. OD is often used as a pre-processing step for downstream tasks such as learning from open-set noisy labels (Wang et al., 2018) and open-set semi-supervised learning (Yu et al., 2020).
2.2 Investigation of Research Activity in Each Field
To investigate the research activity in each field, we comprehensively investigated papers that utilize VLMs from top venues and summarized them in Table 1. In this context, the term “top venues” refers to leading conferences characterized by high impact factors and rigorous peer-review standards, including NeurIPS, AAAI, ICLR, CVPR, ICML, ICCV, ECCV, IJCAI, and ACMMM, together with distinguished journals of comparable impact and recognition, such as TPAMI, IJCV, and TMLR. The authors visited the official pages of each venue and manually counted the number of VLM-based papers by examining their titles and content. Since CLIP, the most representative VLMs, was introduced at ICML 2021, we focused our investigation on the period from 2021 to April 2025. We define the research activity of a field objectively by the number of papers published through a rigorous peer-review process in the period starting with the introduction of VLMs (e.g., CLIP) and continuing to the present.
Our survey reveals that CLIP (Radford et al., 2021) is by far the most widely used VLM for OOD detection, whereas other models such as Grounding DINO (Liu et al., 2024e) and SAM (Kirillov et al., 2023) have seen limited adoption in this context. While many existing methods are developed with CLIP, the concept of OOD detection with CLIP is not limited to this particular model. Therefore, throughout this survey, we use the term VLM-based OOD detection to refer broadly to OOD detection methods that leverage VLMs, and we similarly use the prefix VLM-based for other tasks (e.g., VLM-based AD). However, it is important to note that not all VLMs employ softmax-based scoring mechanisms—for example, SigLIP (Zhai et al., 2023) uses sigmoid-based classification. As a result, some methods (Ming et al., 2022a; Miyai et al., 2025b), which rely on softmax outputs, cannot be directly applied to such models. For clarity and consistency, this survey focuses primarily on CLIP as the representative VLM and investigates VLM-based methods for OOD detection and related tasks within this scope.
As OOD detection research is primarily focused on the image domain, we conduct a survey of other tasks within the image domain that are common and have strong connections to OOD detection research. For instance, our survey does not cover video domain tasks (Du et al., 2024; Zara et al., 2023; Wu et al., 2024b) due to their limited connection to OOD detection. From the results in Table 1, it is clear that Sensory AD and OOD detection have a very high number of papers, while the other fields are not as active. In the next section, we will examine these findings in greater depth and discuss the evolution of each field.
2.3 Generalized OOD Detection v2: Evolution in VLM Era
We propose a generalized OOD detection v2, encapsulating the evolution of each field in the VLM era. The evolution trajectory of the Generalized OOD detection framework v2 is shown in Fig. 2. The evolution of each field is as follows: